EDITORIAL

Terry Goodman

As is our custom, we want to acknowledge those who have served as referees for MJMS for the past year. Their contribution to the journal is so very important and we appreciate the time and energy they give to their task of reviewing submitted manuscripts.

List of Referees

Mangho Ahuja, Cape Girardeau, Missouri Tom Akers, University of Missouri-Rolla Elvan Akin-Bohner, University of Missouri-Rolla Toma Albu, Koc University Eric Aurand, Texas A & M University-Commerce Mohammad K. Azarian, University of Evansville Dale Bachman, Central Missouri State University Margaret Beattie, Mount Allison University Miron Bekker, University of Missouri-Rolla James Benedict, Augusta State College Reginald Brigham, University of Missouri-Rolla Martin Bohner, University of Missouri-Rolla Jim Bruening, Southeast Missouri State University G. Alan Cannon, Southeastern Louisiana University O-Yeat Chan, University of Illinois David M. Clark, SUNY at New Paltz Steve Clark, University of Missouri-Rolla Kumer Das, Auburn University Joseph B. Dence, University of Missouri-St. Louis Thomas Dence, Ashland University David Drain, University of Missouri-Rolla Roman Dwilewicz, University of Missouri-Rolla Keith Emmert, Texas Tech University Russell Euler, Northwest Missouri State University Joe Flowers, San Antonio, Texas Gary Gadbury, University of Missouri-Rolla Ramesh Garimella, University of Central Arkansas

List of Referees (cont.)

Louis Grimm, University of Missouri-Rolla Paul Hill, Western Kentucky University Joe Howard, New Mexico Highlands University Eric Howington, Coastal Carolina University Matt Insall, University of Missouri-Rolla Robert E. Kennedy, Pleasant Hill, Missouri Vlajko Kocic, Xavier University of Louisiana Alessandro Languasco, University of Padova Edward S. Letzter, Temple University Wayne McDaniel, University of Missouri-St. Louis Timothy McNicholl, Lamar University Ronald Mickens, Clark Atlanta University Kent M. Neuerburg, Southeastern Louisiana University James P. Ochoa, Hardin-Simmons University Les Reid, Southwest Missouri State University Robert Roe, University of Missouri-Rolla Paul Runnion, University of Missouri-Rolla Mayumi Sakata, William Jewell College Suman Sanyal, University of Missouri-Rolla Philip Schultz, University of Western Australia James Sly, University of Missouri-Rolla Allen Stenger, Alamogordo, New Mexico Songlin Tian, Central Missouri State University Peter Trotter, University of Tasmania Jerry Veeh, Auburn University Donald Vestal, Missouri Western State University Hristo Voulov, University of Missouri-Kansas City Haohao Wang, Southeast Missouri State University Lianwen Wang, Central Missouri State University Howard Warth, University of Missouri-Rolla Lawrence Washington, University of Maryland David Weinreich, Arlington, VA Christina Welch, University of Missouri-Rolla Mahmoud Yousef, Central Missouri State University